OpenStudy (anonymous):

How do I start to get the integral of dx/1+sinx ? Do I multiply with 1-sinx/ 1-sinx ?

hartnn (hartnn):

do u know this standard substitution? t= tan (x/2) ?

OpenStudy (anonymous):

It is in the the textbook, but I really struggle with it.

hartnn (hartnn):

can u find sin x in terms of t with t= tan (x/2) ??

hartnn (hartnn):

or do u know the formula of sin 2x in terms of tan x ??

hartnn (hartnn):

let me tell u sin2x = 2 tan x /(1+ tan^2 x) so sin x = 2tan (x/2)/(1+ tan ^2 (x/2)) = 2t/(1+t^2) and since x= 2 tan^-1 t dx = 2/(1+t^2) clear till here ?

OpenStudy (lgbasallote):

i think @Iventer 's method will work too...

OpenStudy (anonymous):

i want to write it down to check it. i only have sin2x = 2sinxcosx

OpenStudy (anonymous):

how do you get to the tan formula

hartnn (hartnn):

its standard , remember it. and yes , inverter's way will also work, but i want inverter to know this substitution as it helps in many different problems.

OpenStudy (lgbasallote):

i've never used half-angle theorems in integrals though

OpenStudy (lgbasallote):

all i know is the formula for \[\sin (\alpha \pm \beta)\] \[\cos (\alpha \pm \beta)\] then the pythagorean identities and it has kept me happy for a long time

hartnn (hartnn):

if t= tan (x/2) cos x = (1-t^2)/(1+t^2) sin x = (2t)/(1+t^2) dx = (2)/ ( 1+t^2) note it down.

OpenStudy (lgbasallote):

i suggest you memorize as little as possible @Iventer and just practice and learn to manipulate your surroundings to your advantage

OpenStudy (lgbasallote):

im not saying dont listen to hartnn though

OpenStudy (anonymous):

ok i have that @hartnn

hartnn (hartnn):

now substitute dx and sin x in your integral , you should get 2/((1+t)^2)

OpenStudy (anonymous):

give me second

OpenStudy (anonymous):

ok i have it. what next

hartnn (hartnn):

can u integrate that ?

hartnn (hartnn):

what answer u have in your answer key?

hartnn (hartnn):

is it in half angle form ??

hartnn (hartnn):

else this entire procedure will be more difficult than to multiply and divide by (1-sin x)

OpenStudy (anonymous):

we dont really work with the half angle. i am really lost as i am studying long distance and have to figure all this out on my own with just a textbook as help. I have to know the tan(x/2) method as well, but i think i will have to look at it more. I can use what u have given me. but i need more time. i will close for now but get back later. thank you so much for your help so far xxxxx

hartnn (hartnn):

ok,going with your original question,YES, multiply and divide by (1-sin x) then u get (1-sin^2 x) = cos^2 x in the denominator, then just separate out the denominators, and u will get 1/(cos ^2 x) - sinx/ cos^2 x = sec^2 x - sec x tan x now this is easy to integrate, isn't it ?

OpenStudy (anonymous):

much better, yes thank u lots

hartnn (hartnn):

there are many approaches to solve a particular integral, here it was simpler to use (1-sin x). But in problems like integral of {1/(a sinx + b cos x +c)} dx, that substitution will be simplest to solve....maybe in near future, u have to deal with such integrals.

OpenStudy (anonymous):

i have one question that is asking to solve that way, so yes i have to look at it in near future. but i have idea where to start now with sin, cos and dx u gave me. much appreciated